OpenStudy (anonymous):

Pedro owns 24 book. If 1/3 are science and 1/4 are history, how many of his books are about other topics?

OpenStudy (anonymous):

just multiply 24 books by 1/3 = 8 books then 24 books by 1/4 = 6 books. So 8 + 6 = 14. Subtract the 14 books from 24 to get the remaining book of other topics..

OpenStudy (anonymous):

Thank you!!! you were very helpful.

OpenStudy (anonymous):

your welcome